# Asymmetric Errors

### v1

@@ A.8 The Edgeworth expansion
The Edgeworth expansion [Edgeworth] is appropriate with an asymmetry ascribable to the fact that the sample size is finite and the Central Limit Theorem has not had a chance to make the distribution accurately Gaussian. This is not a probability transform so there is no equivalent to Figures 18 or 20. The pdf (taking only the first term after the Gaussian) is given by
$p(x)={1\over\sigma\sqrt{2\pi}}e^{-{1\over 2}(x-\mu)^{2}/\sigma^{2}}\left(1+{\gamma\over 6\sigma^{3}}He_{3}({x-\mu\over\sigma})\right)$ (40)
@@
$F(z)=\Phi(z)-{\gamma\over 6\sigma^{3}}\phi(z)(z^{2}-1)$ (41)
where $z\equiv(x-\mu)/\sigma$ and $\phi(z)$ and $\Phi(z)$ are the Gaussian density and distribution functions. The second term vanishes at the 1-sigma points, $z=\pm 1$ so the width of the 68% central interval is the same for all values of $\gamma$. As $\mu$ is also independent of $\gamma$, this means that the choice of parameters ${\mu}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ cannot be used with this form, as mentioned in Section 2.1.
@@
$\mu$ $=$ $M+{1\over 2}({\sigma^{+}}-{\sigma^{-}}),$
$\sqrt{V}\equiv\sigma$ $=$ ${1\over 2}({\sigma^{+}}+{\sigma^{-}}),$ (42)
@@
$z=\Phi^{-1}\left({1\over 2}+{\gamma\over 6\sigma^{3}}(z^{2}-1)G(z)\right)$ (43)
@@ A.9 The Skew Normal
Azzalini [azzalini] suggests using a distribution whose density can be elegantly written as
@@
In addition to the asymmetry parameter $\alpha$ one introduces a scale parameter $\omega$ and a location parameter $\xi$, so $z\equiv{x-\xi\over\omega}$. Writing for convenience $\delta={\alpha\over\sqrt{1+\alpha^{2}}}$ the moments are given by
@@
$\gamma$ $=$ $\left({4-\pi\over 2}\right)(\omega\delta\sqrt{2/\pi})^{3}$
@@
$\xi$ $=$ $\mu-\omega\delta\sqrt{2\over\pi}$
@@
To convert in the other direction, when the distribution is specified by ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ the parameter $\alpha$ can be found numerically by adjusting it to get the desired value of the asymmetry $A={{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}$. $\omega$ is then found by scaling to get the desired value of ${\sigma^{+}}+{\sigma^{-}}$, and $\xi$ from the desired value for $M$. If the distribution is specified by ${\mu}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ then to obtain the other forms is not impossible (as it is for the Edgeworth distribution) but is so impractical as to be essentially useless.
@@ A.10 The Johnson system
The combination of Johnson distributions $S_{B}$ (bounded), $S_{U}$ (unbounded), together with their limiting cases (Gaussian and log-normal), form a very flexible system of four-parameter densities [ref:johnson, ref:johnsonbook, ref:hahnshap]. It turns out that there is a Johnson distribution for every possible combination of mean, standard deviation, skewness, and kurtosis. It is in fact convenient to think of the set as if it was a single distribution whose density is given by $j(R,\mu,\sigma,s,\kappa)$ (in this parameterization, skewness $s$ and kurtosis $\kappa$ are normalized).
@@
$\rho(x)=\frac{\delta}{\lambda\sqrt{2\pi\left(1+\left(\frac{x-\xi}{\lambda}\right)^{2}\right)}}\,e^{-\frac{1}{2}\left(\gamma+\delta\,\mbox{\scriptsize sinh}^{-1}\left(\frac{x-\xi}{\lambda}\right)\right)^{2}},$ (49)
@@
With the notation $\Omega=\gamma/\delta$ and $\omega=\exp(\delta^{-2})$, the moments of the $S_{U}$ distribution are given by [ref:johnsonbook]
$\begin{split}\mu&=\xi-\lambda\sqrt{\omega}\,\mbox{sinh}\,\Omega\\ \sigma^{2}&=\frac{\lambda^{2}}{2}(\omega-1)(\omega\,\mbox{cosh}\,2\Omega+1)\\ s^{2}&=\frac{\omega(\omega-1)\left[\omega(\omega+2)\,\mbox{sinh}\,3\Omega+3\,\mbox{sinh}\,\Omega\right]^{2}}{2(\omega\,\mbox{cosh}\,2\Omega+1)^{3}}\\ \kappa&=\frac{\omega^{2}(\omega^{4}+2\omega^{3}+3\omega^{2}-3)\,\mbox{cosh}\,4\Omega+4\omega^{2}(\omega+2)\,\mbox{cosh}\,2\Omega+3(2\omega+1)}{2(\omega\,\mbox{cosh}\,2\Omega+1)^{2}}\end{split}$ (51)
The sign of $s$ is opposite to that of $\gamma$. The mapping from the parameter set $\{\mu,\sigma,s,\kappa\}$ back to $\{\xi,\lambda,\gamma,\delta\}$ can be efficiently performed numerically, according to the algorithm presented in [ref:AS99]. The calculation of $\hat{\kappa}$ for the given $s$ also has to be performed numerically, as well as the determination of $\mu$, $\sigma$, and $s$ from quantiles.
An example $S_{U}$ density and the corresponding transformation are shown in Fig. 26.
Figure 26: The panel on the right illustrates the Johnson $S_{U}$ density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Eq. 50. As expected, this transformation maps $-1$ to $-0.6$, 00 to 00, and $1$ into $1$.
A.11 The Log-normal
The log-normal distribution is a special three-parameter member of the Johnson system. For a fixed skewness, the kurtosis of log-normal is larger than the kurtosis of all $S_{B}$ distributions but smaller than the kurtosis of all $S_{U}$. It is thus the limiting distribution for $S_{U}$ with the smallest possible kurtosis and for $S_{B}$ with the largest possible one.
A variable distributed according to the log-normal can be obtained from the standard normal variate $z$ by applying the transformation
@@ A.11 The Log-normal
This also means that the variable $z=\gamma+\delta\ln(x-\xi)$ is distributed according to the standard normal if $x$ is distributed according to the log-normal. With this parameterization [^2] The parameterization described here is consistent with [ref:johnsonbook] but differs from the standard one., for positive values of skewness the log-normal density is given by
$\rho(x)=\begin{cases}0,&x\leq\xi\\ \frac{\delta}{\sqrt{2\pi}(x-\xi)}e^{-\frac{1}{2}\left[\gamma+\delta\ln(x-\xi)\right]^{2}},&x>\xi\end{cases}$ (53)
Distributions with negative values of skewness can be obtained by utilizing $\rho(-x)$ and adjusting the parameters $\xi$ and $\gamma$ as necessary.
With the notation $\omega=\exp(\delta^{-2})$, the moments of the log-normal distribution (with normalized $s$ and $\kappa$) are given by [ref:johnsonbook]
$\begin{split}\mu&=\xi+\sqrt{\omega}\exp\left(-\frac{\gamma}{\delta}\right)\\ \sigma^{2}&=\omega(\omega-1)\exp\left(-\frac{2\gamma}{\delta}\right)\\ s^{2}&=(\omega-1)(\omega+2)^{2}\\ \kappa&=\omega^{4}+2\omega^{3}+3\omega^{2}-3\end{split}$ (54)
The mapping from the parameter set $\{\mu,\sigma,s\}$ back to $\{\xi,\gamma,\delta\}$ can be performed by first solving the equation $s^{2}=(\omega-1)(\omega+2)^{2}$ for $\omega$, calculating $\delta$, and then finding all other parameters. See [ref:johnsonbook] for details. The mapping from the quantiles to the distribution parameters has to be constructed numerically.
An example log-normal density and the corresponding transformation are shown in Fig. 27.
Figure 27: The panel on the right illustrates the log-normal density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Eq. 52.
@@ Appendix B Modelling likelihood errors
There are many other possible forms that could be used to approximate a parabola (for example, the ‘rotated parabola’) or then by taking the logarithm of the pdf models described in Appendix A. The software provided is flexible so that more models (including models with more parameters) can be added without needing to re-write all the code.
@@ B.1 The cubic
$\alpha={{\sigma^{+}}^{3}+{\sigma^{-}}^{3}\over{\sigma^{+}}^{2}{\sigma^{-}}^{2}({\sigma^{+}}+{\sigma^{-}})}\qquad\beta={{\sigma^{-}}^{2}-{\sigma^{+}}^{2}\over{\sigma^{+}}^{2}{\sigma^{-}}^{2}({\sigma^{+}}+{\sigma^{-}})}.$ (56)
This gives curves which behave sensibly in the range $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$ if ${\sigma^{+}}$ and ${\sigma^{-}}$ differ by less than a factor of 2 (if the asymmetry is large, the curve will have a minimum inside the range in addition to the maximum at $\hat{a}=0$), but outside that range the cubic term gives an unwanted turning point, and the curve does not go to $\infty$- ∞ at both large positive and negative $a$.
@@ B.2 The broken parabola
$\ln L(a)=g(a)=\begin{cases}-\frac{1}{2}\frac{a^{2}}{{\sigma^{-}}^{2}}&\text{for }a\leq 0\\ -\frac{1}{2}\frac{a^{2}}{{\sigma^{+}}^{2}}&\text{for }a\geq 0\\ \end{cases}$ (57)
(here, we explicitly assume $\hat{a}=0$). Unfortunately, this solution suffers from the unphysical discontinuity of the second derivative at $a=0$.
@@ B.3 The symmetrized parabola
The symmetrized parabola model is constructed in a Bayesian manner. It starts by assuming a flat prior for the parameter and then treats Eq. 57 as a logarithm of the (unnormalized) parameter probability density. The resulting distribution is Fechner with density given by Eq. 38 in which $m=\hat{a}$, $\sigma_{1}={\sigma^{-}}$, and $\sigma_{2}={\sigma^{+}}$. This distribution is then exchanged for the Gaussian distribution with the same mean and standard deviation as the Fechner (the moments are calculated according to Eq. 39). The symmetrized parabola is the logarithm of this Gaussian density.
Note that the peak position of the symmetrized parabola is no longer at $\hat{a}$, it is shifted to the right if ${\sigma^{+}}>{\sigma^{-}}$ or to the left in case ${\sigma^{+}}<{\sigma^{-}}$.
@@ B.4 The constrained quartic
$(\ln L)^{\prime\prime}(a)=-{1\over 2}(\alpha+\beta a)^{2}\qquad\ln L(a)=-{1\over 2}\left({\alpha^{2}a^{2}\over 2}+{\alpha\beta a^{3}\over 3}+{\beta^{2}a^{4}\over 12}\right)$ (58)
@@
The $\pm$ term should be taken as negative to give a small quartic term. In very asymmetric cases (${\sigma^{+}}$ and ${\sigma^{-}}$ differing by more than a factor of $(1+\sqrt[4]{12}+\sqrt{3})/2\approx 2.29663$) the inner root is negative, indicating that there is no solution in the desired form.
@@
This gives better behaviour than the cubic for large $a$, but is not always so satisfactory in the central $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$ region
@@ B.5 The molded quartic
$\begin{split}\alpha&=3({\sigma^{-}}-{\sigma^{+}})^{2}(5{\sigma^{-}}^{6}+8{\sigma^{-}}^{5}{\sigma^{+}}+5{\sigma^{-}}^{4}{\sigma^{+}}^{2}+8{\sigma^{-}}^{3}{\sigma^{+}}^{3}+5{\sigma^{-}}^{2}{\sigma^{+}}^{4}+8{\sigma^{-}}{\sigma^{+}}^{5}+5{\sigma^{+}}^{6})/\eta\\ \beta&=({\sigma^{-}}-{\sigma^{+}})[25({\sigma^{-}}^{8}+{\sigma^{+}}^{8})+14({\sigma^{-}}^{7}{\sigma^{+}}-{\sigma^{-}}^{6}{\sigma^{+}}^{2}+{\sigma^{-}}^{5}{\sigma^{+}}^{3}-{\sigma^{-}}^{4}{\sigma^{+}}^{4}+{\sigma^{-}}^{3}{\sigma^{+}}^{5}-{\sigma^{-}}^{2}{\sigma^{+}}^{6}+{\sigma^{-}}{\sigma^{+}}^{7})]/\eta\\ \gamma&=(10{\sigma^{-}}^{10}-5{\sigma^{-}}^{9}{\sigma^{+}}+30{\sigma^{-}}^{7}{\sigma^{+}}^{3}-6{\sigma^{-}}^{6}{\sigma^{+}}^{4}+6{\sigma^{-}}^{5}{\sigma^{+}}^{5}-6{\sigma^{-}}^{4}{\sigma^{+}}^{6}+30{\sigma^{-}}^{3}{\sigma^{+}}^{7}-5{\sigma^{-}}{\sigma^{+}}^{9}+10{\sigma^{+}}^{10})/\eta\end{split}$ (62)

### v2

@@ A.8 The Edgeworth expansion
The Edgeworth expansion [19] is appropriate with an asymmetry ascribable to the fact that the sample size is finite and the Central Limit Theorem has not had a chance to make the distribution accurately Gaussian. This is not a probability transform so there is no equivalent to Figures 18 or 20. The pdf (taking only the first term after the Gaussian) is given by
$P(x)={1\over\sigma\sqrt{2\pi}}e^{-{1\over 2}(x-\mu)^{2}/\sigma^{2}}\left[1+{\gamma\over 6\sigma^{3}}He_{3}({x-\mu\over\sigma})\right],$ (38)
@@
$F(z)=\Phi(z)-{\gamma\over 6\sigma^{3}}\phi(z)(z^{2}-1),$ (39)
where $z\equiv(x-\mu)/\sigma$ and $\phi(z)$ and $\Phi(z)$ are the Gaussian density and distribution functions. The second term vanishes at the 1-sigma points, $z=\pm 1$, so the width of the 68% central interval is the same for all values of $\gamma$. As $\mu$ is also independent of $\gamma$, this means that the choice of parameters ${\mu}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ cannot be used with this form, as mentioned in Section 2.1.
@@
$\mu$ $=$ $M+{1\over 2}({\sigma^{+}}-{\sigma^{-}})$
$\sqrt{V}\equiv\sigma$ $=$ ${1\over 2}({\sigma^{+}}+{\sigma^{-}})$ (40)
@@
$z=\Phi^{-1}\left[{1\over 2}+{\gamma\over 6\sigma^{3}}(z^{2}-1)G(z)\right],$ (41)
@@ A.9 The skew normal
Azzalini [22] suggests using a distribution whose density can be elegantly written as
@@
In addition to the asymmetry parameter $\alpha$ one introduces a scale parameter $\omega$ and a location parameter $\xi$, so $z\equiv{x-\xi\over\omega}$. Writing for convenience $\delta={\alpha\over\sqrt{1+\alpha^{2}}}$, the moments are given by
@@
$\gamma$ $=$ $\left({4-\pi\over 2}\right)(\omega\delta\sqrt{2/\pi})^{3}.$
@@
$\xi$ $=$ $\mu-\omega\delta\sqrt{2\over\pi}.$
@@
To convert in the other direction, when the distribution is specified by ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$, the parameter $\alpha$ can be found numerically by adjusting it to get the desired value of the asymmetry ${{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}$. $\omega$ is then found by scaling to get the desired value of ${\sigma^{+}}+{\sigma^{-}}$, and $\xi$ from the desired value for $M$. If the distribution is specified by ${\mu}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ then to obtain the other forms is not impossible (as it is for the Edgeworth distribution), but is so impractical as to be essentially useless.
@@ A.10 The Johnson system
The combination of Johnson distributions $S_{B}$ (bounded), $S_{U}$ (unbounded), together with their limiting cases (Gaussian and log normal), form a very flexible system of four-parameter densities [23, 24, 25]. It turns out that there is a Johnson distribution for every possible combination of mean, standard deviation, skewness, and kurtosis. It is in fact convenient to think of the set as if it was a single distribution whose density is given by $j(R,\mu,\sigma,s,\kappa)$. (In this parameterization, skewness $s$ and kurtosis $\kappa$ are normalized).
@@
$P(x)=\frac{\delta}{\lambda\sqrt{2\pi\left[1+\left(\frac{x-\xi}{\lambda}\right)^{2}\right]}}\,e^{-\frac{1}{2}\left[\gamma+\delta\,\mbox{\scriptsize sinh}^{-1}\left(\frac{x-\xi}{\lambda}\right)\right]^{2}},$ (47)
@@
With the notation $\Omega=\gamma/\delta$ and $\omega=\exp(\delta^{-2})$, the moments of the $S_{U}$ distribution are given by [24]
$\begin{split}\mu&=\xi-\lambda\sqrt{\omega}\,\mbox{sinh}\,\Omega\\ \sigma^{2}&=\frac{\lambda^{2}}{2}(\omega-1)(\omega\,\mbox{cosh}\,2\Omega+1)\\ s^{2}&=\frac{\omega(\omega-1)\left[\omega(\omega+2)\,\mbox{sinh}\,3\Omega+3\,\mbox{sinh}\,\Omega\right]^{2}}{2(\omega\,\mbox{cosh}\,2\Omega+1)^{3}}\\ \kappa&=\frac{\omega^{2}(\omega^{4}+2\omega^{3}+3\omega^{2}-3)\,\mbox{cosh}\,4\Omega+4\omega^{2}(\omega+2)\,\mbox{cosh}\,2\Omega+3(2\omega+1)}{2(\omega\,\mbox{cosh}\,2\Omega+1)^{2}}.\end{split}$ (49)
The sign of the skewness $s$ is opposite to that of $\gamma$. The mapping from the parameter set $\{\mu,\sigma,s,\kappa\}$ back to $\{\xi,\lambda,\gamma,\delta\}$ can be efficiently performed numerically, according to the algorithm presented in [26]. The calculation of $\hat{\kappa}$ for the given $s$ also has to be performed numerically, as well as the determination of $\mu$, $\sigma$, and $s$ from quantiles.
An example $S_{U}$ density and the corresponding transformation are shown in Figure 26.
Figure 26: The Johnson distribution: the panel on the right illustrates the Johnson $S_{U}$ density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 48. As expected, this transformation maps $-1$ to $-0.6$, 00 to 00, and $1$ into $1$.
A.11 The log normal
The log normal distribution is a special three-parameter member of the Johnson system. For a fixed skewness, the kurtosis of the log normal is larger than the kurtosis of all $S_{B}$ distributions, but smaller than the kurtosis of all $S_{U}$. It is thus the limiting distribution for $S_{U}$ with the smallest possible kurtosis and for $S_{B}$ with the largest possible one.
A variable distributed according to the log normal can be obtained from the standard normal variate $z$ by applying the transformation
@@ A.11 The log normal
This also means that the variable $z=\gamma+\delta\ln(x-\xi)$ is distributed according to the standard normal if $x$ is distributed according to the log normal. With this parameterization [^3] The parameterization described here is consistent with [24], but differs from the standard one., for positive values of skewness, the log normal density is given by
$P(x)=\begin{cases}0,&x\leq\xi\\ \frac{\delta}{\sqrt{2\pi}(x-\xi)}e^{-\frac{1}{2}\left[\gamma+\delta\ln(x-\xi)\right]^{2}},&x>\xi\end{cases}.$ (51)
Distributions with negative values of skewness can be obtained by utilizing $P(-x)$ and adjusting the parameters $\xi$ and $\gamma$ as necessary.
With the notation $\omega=\exp(\delta^{-2})$, the moments of the log normal distribution (with normalized $s$ and $\kappa$) are given by [24]
$\begin{split}\mu&=\xi+\sqrt{\omega}\exp\left(-\frac{\gamma}{\delta}\right)\\ \sigma^{2}&=\omega(\omega-1)\exp\left(-\frac{2\gamma}{\delta}\right)\\ s^{2}&=(\omega-1)(\omega+2)^{2}\\ \kappa&=\omega^{4}+2\omega^{3}+3\omega^{2}-3\end{split}.$ (52)
The mapping from the parameter set $\{\mu,\sigma,s\}$ back to $\{\xi,\gamma,\delta\}$ can be performed by first solving the equation $s^{2}=(\omega-1)(\omega+2)^{2}$ for $\omega$, calculating $\delta$, and then finding all other parameters. See [24] for details. The mapping from the quantiles to the distribution parameters has to be constructed numerically.
An example log normal density and the corresponding transformation are shown in Fig. 27.
Figure 27: The log normal distribution: the panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 50.
@@ Appendix B Modelling likelihood errors
Here we describe several models for likelihood curves that are approximately parabolic. There are many other possible forms (for example, the ‘rotated parabola’), and both implementations of software provided are flexible so that more models (including models with more parameters) can be added without needing to re-write all the code.
@@ B.1 The cubic
$\alpha={{\sigma^{+}}^{3}+{\sigma^{-}}^{3}\over{\sigma^{+}}^{2}{\sigma^{-}}^{2}({\sigma^{+}}+{\sigma^{-}})}\qquad{\rm and}\qquad\beta={{\sigma^{-}}^{2}-{\sigma^{+}}^{2}\over{\sigma^{+}}^{2}{\sigma^{-}}^{2}({\sigma^{+}}+{\sigma^{-}})}.$ (54)
If ${\sigma^{+}}$ and ${\sigma^{-}}$ differ by less than a factor of 2 this gives curves which behave sensibly in the range $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$, but outside that range the cubic term gives an unwanted turning point. (If the asymmetry is large, the curve will have a minimum inside the range in addition to the maximum at $\hat{a}=0$.) Also the curve does not go to $\infty$- ∞ at both large positive and negative $a$. So although the cubic is the obvious extension of the quadratic, it often gives problems.
@@ B.2 The broken parabola
$\ln L(a)=g(a)=\begin{cases}-\frac{1}{2}\frac{a^{2}}{{\sigma^{-}}^{2}}&\text{for }a\leq 0\\ -\frac{1}{2}\frac{a^{2}}{{\sigma^{+}}^{2}}&\text{for }a\geq 0\\ \end{cases}.$ (55)
(Here, we explicitly assume $\hat{a}=0$.) Unfortunately, this solution suffers from the unphysical discontinuity of the second derivative at $a=0$.
@@ B.3 The symmetrized parabola
The symmetrized parabola model is constructed in a Bayesian manner. It starts by assuming a flat prior for the parameter and then treats Equation 55 as a logarithm of the (unnormalized) parameter probability density. The resulting distribution is Fechner with density given by Equation 36 in which $m=\hat{a}$, $\sigma_{1}={\sigma^{-}}$, and $\sigma_{2}={\sigma^{+}}$. This distribution is then exchanged for the Gaussian distribution with the same mean and standard deviation as the Fechner where the moments are calculated according to Equation 37. The symmetrized parabola is the logarithm of this Gaussian density.
Note that the peak position of the symmetrized parabola is no longer at $\hat{a}$; it is shifted to the right if ${\sigma^{+}}>{\sigma^{-}}$ or to the left in case ${\sigma^{+}}<{\sigma^{-}}$.
@@ B.4 The constrained quartic
$(\ln L)^{\prime\prime}(a)=-{1\over 2}(\alpha+\beta a)^{2}\qquad{\rm which\ yields}\qquad\ln L(a)=-{1\over 2}\left({\alpha^{2}a^{2}\over 2}+{\alpha\beta a^{3}\over 3}+{\beta^{2}a^{4}\over 12}\right)$ (56)
@@
The $\pm$ term should be taken as negative to give a small quartic term. In very asymmetric cases, where ${\sigma^{+}}$ and ${\sigma^{-}}$ differ by more than a factor of $(1+\sqrt[4]{12}+\sqrt{3})/2\approx 2.29663$, the inner root is negative, indicating that there is no solution in the desired form.
@@
This gives better behaviour than the cubic for large $a$, but is not always so satisfactory in the central $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$ region.
@@ B.5 The molded quartic
$\begin{split}\alpha&=3({\sigma^{-}}-{\sigma^{+}})^{2}(5{\sigma^{-}}^{6}+8{\sigma^{-}}^{5}{\sigma^{+}}+5{\sigma^{-}}^{4}{\sigma^{+}}^{2}+8{\sigma^{-}}^{3}{\sigma^{+}}^{3}+5{\sigma^{-}}^{2}{\sigma^{+}}^{4}+8{\sigma^{-}}{\sigma^{+}}^{5}+5{\sigma^{+}}^{6})/\eta\\ \beta&=({\sigma^{-}}-{\sigma^{+}})[25({\sigma^{-}}^{8}+{\sigma^{+}}^{8})+14({\sigma^{-}}^{7}{\sigma^{+}}-{\sigma^{-}}^{6}{\sigma^{+}}^{2}+{\sigma^{-}}^{5}{\sigma^{+}}^{3}-{\sigma^{-}}^{4}{\sigma^{+}}^{4}+{\sigma^{-}}^{3}{\sigma^{+}}^{5}-{\sigma^{-}}^{2}{\sigma^{+}}^{6}+{\sigma^{-}}{\sigma^{+}}^{7})]/\eta\\ \gamma&=(10{\sigma^{-}}^{10}-5{\sigma^{-}}^{9}{\sigma^{+}}+30{\sigma^{-}}^{7}{\sigma^{+}}^{3}-6{\sigma^{-}}^{6}{\sigma^{+}}^{4}+6{\sigma^{-}}^{5}{\sigma^{+}}^{5}-6{\sigma^{-}}^{4}{\sigma^{+}}^{6}+30{\sigma^{-}}^{3}{\sigma^{+}}^{7}-5{\sigma^{-}}{\sigma^{+}}^{9}+10{\sigma^{+}}^{10})/\eta\end{split}.$ (60)